OpenStudy (anonymous):

find a third-degree polynomial function such as f(0)=18 & whos zeros are -1,2,& 3. using complete sentences explain how you found it

OpenStudy (anonymous):

f(x)=a(x+1)(x-2)(x-3) to find a replace f(0)=18

OpenStudy (anonymous):

what's next after that?

OpenStudy (whpalmer4):

You have to expand the polynomial, set it = to 18 and solve for the value of \(a\).

OpenStudy (whpalmer4):

(at x = 0)

OpenStudy (anonymous):

would be 18=a(6) ?

OpenStudy (whpalmer4):

\[f(0) = 18 = a(0+1)(0-2)(0-3) = 6a \]\[a =\]

OpenStudy (whpalmer4):

Probably you need to multiply the polynomial out rather than leaving it in factored form. Even if you don't need to, you should — most people here need the practice! (and if you don't need the practice, you'll be able to do it with no effort)